# Fast Revenue Maximization

### v1

@@ 5 Auxiliary Results
Proof of Lemma 1.
The single crossing result for regular distributions mentioned in Lemma 1 follows directly from (Allouah et al., 2022, Lemma 2) with the following additional identity:
For General distributions, the result follows directly from the non-increasing property of CCDF functions. ∎
@@
For any $r^{*}$ such that ${\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-empty, $\sup{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})=\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
@@
Step 1: We first show that, $\sup{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\leq\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Let $q\in{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ and let $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ such that ${r^{*}}$ is an optimal price for $F$ with associated conversion rate $q$. Consider $i\in\{0,\ldots,N\}$ such that ${r^{*}}\in[p_{i},p_{i+1}]$ The crossing lemma implies that, $q=\overline{F}({r^{*}})\leq\min\left(\overline{G}_{\pazocal{F},{i-1}}({r^{*}}),\overline{G}_{\pazocal{F},{i+1}}({r^{*}})\right)\leq\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}}),$ where the last inequality holds by definition of $\overline{U}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ when $\pazocal{C}$ is either $\pazocal{G}$ or $\pazocal{F}$.
@@
Step 2: Next, we show that $\sup{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\geq\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ by proving that $\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. To do so, we need to show that there exists a distribution $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ such that ${r^{*}}$ is an optimal price and $\overline{F}({r^{*}})=\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})$. We consider as a candidate the distribution $F_{\pazocal{C}}(\cdot|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}})$ defined in Equation 2.10. Let $\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+}=\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}}))$ and note that, $U_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$ and therefore this set is non-empty. Hence, by Lemma 3, $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$, which implies that, $F_{\pazocal{C}}(\cdot|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}})\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\subset\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
@@
To show that ${r^{*}}$ is the optimal price, we see from Lemma 4 that it suffices to show that ${r^{*}}\cdot\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})\geq\max_{i}p_{i}\cdot q_{i}$. Let $q\in{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ then by definition there exists $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ such that $\overline{F}({r^{*}})=q$ and ${r^{*}}\cdot\overline{F}({r^{*}})\geq p\cdot\overline{F}(p)$ for every $p\in[\underline{v},\overline{v}]$. This implies that for every $i\in\{0,N+1\}$,
${r^{*}}\cdot\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})\stackrel{{\scriptstyle(a)}}{{\geq}}{r^{*}}\cdot q={r^{*}}\cdot\overline{F}({r^{*}})\geq p_{i}\cdot\overline{F}(p_{i})=p_{i}\cdot q_{i},$
@@
Appendix for
Fast revenue maximization with few experiments
Achraf Bahamou, Columbia University, IEOR department, achraf.bahamou@columbia.edu
Omar Besbes, Columbia University, Graduate School of Business, obesbes@columbia.edu
Omar Mouchtaki, NYU Stern School of Business, om2166@nyu.edu
Appendix A Details on numerical evaluations in Section 3
In Section 3, we provide an understanding of the best achievable Ratio across all pricing mechanisms in $\pazocal{P}$ for different choices of $\mathscr{I}_{\mathbf{p},\mathbf{q}}$.
We next, provide more details about how we numerically derive a lower bound on (Maximin) by leveraging the reduction in Theorem 1. First, we note that the epigraph formulation of (Maximin) satisfies,
$\!\sup_{\Psi(\cdot)\in\pazocal{P},\,\lambda\in[0,1]}$ $\lambda$
$\!\sup_{\Psi(\cdot)\in\pazocal{P},\,\lambda\in[0,1]}$ $\lambda$
subject to $R_{\lambda}\left(\Psi,F_{\pazocal{C}}(\cdot|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)\leq 0\quad\text{ for every }{r^{*}}\in[\underline{v},\overline{v}]\text{ s.t. }{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\neq\emptyset.$
Finally, we derive a lower bound on Problem (Maximin) by restricting the space of randomized mechanisms to discrete mechanisms (i.e., mechanisms such that the distribution of prices a discrete over $[\underline{v},\overline{v}]$) and projecting the one dimensional set of constraints into the associated support of the mechanisms. More specifically, consider an increasing sequence of $M+2$ positive reals $\mathbb{A}=\{a_{i}\}_{i=0}^{M+1}$ a partition of the feasible interval $\mathscr{S}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ for $M\mbox{ in }\mathbb{N}^{*}$ and define the set of discrete mechanisms on $\mathbb{A}$ as
$\pazocal{P}_{\mathbb{A}}=\left\{\Psi\in\pazocal{P}\>:\>\Psi(x)=\sum_{j=0}^{M+1}p_{j}\mathbf{1}\{x\geq a_{i}\},\mbox{ for some }0\leq p_{i}\leq 1,\>\sum_{j=0}^{M+1}p_{i}=1\right\}.$
$\underline{\pazocal{L}}(\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}),\mathbb{A})=$ $\!\sup_{\mathbf{p},\lambda}$ $\lambda$
subject to $\lambda\cdot\mbox{opt}\left(\overline{F}_{\pazocal{C}}(\cdot|a_{i+1}^{-},\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)-\sum_{j=0}^{M}Rev\left(a_{j}|\overline{F}_{\pazocal{C}}(\cdot|a_{i},\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)p_{j}\leq 0\quad i=0,...M+1,$
$\sum_{j=0}^{M+1}p_{j}\leq 1,\quad p_{i}\geq 0\quad i=0,...M+1,\quad 0\leq\lambda\leq 1,$
with $\mbox{opt}(F_{\pazocal{C}}(\cdot|a_{i+1}^{-},\mathscr{I}_{\mathbf{p},\mathbf{q}}))=\lim_{x\to a_{i+1}^{-}}\mbox{opt}(F_{\pazocal{C}}(\cdot|x,\mathscr{I}_{\mathbf{p},\mathbf{q}}))$ for any $i=0,...M$.
We note that a lines of technical arguments similar to the one used in (Allouah et al., 2023) enable to show that $\underline{\pazocal{L}}(\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}),\mathbb{A})$ is a lower bound on the value of (Maximin). Furthermore, the solution $(p_{i})_{i\leq N+2}$ provides a feasible distribution of prices. Finally, we note that by judiciously constructing a discrete grid one can ensure that the values associated with the sequence of linear programs converge to the one of (Maximin) as the grid becomes finer.
In all the applications discussed in Section 3, we use the grid of $M=2500$.
Appendix B Details about Section 3.3
We first present formally the Ternary Search procedure in Algorithm 2.
Initialize $a=\underline{v}$, $b=\overline{v}$, $\pazocal{H}_{0}=\{(\underline{v},1),(\overline{v},0)\}$;
for t=1,2,3… do
       Price at $p_{1}=a+\frac{b-a}{3}$, observe $q_{1}$ ;
       Price at $p_{2}=a+\frac{2(b-a)}{3}$, observe $q_{2}$ ;
       Record observed conversion rates $\pazocal{H}_{t+1}=\pazocal{H}_{t}\cup\{(p_{1},q_{1}),(p_{2},q_{2})\}$ ;
      if $p_{1}\cdot q_{1}<p_{2}\cdot q_{2}$ then
             Set $a=p_{1}$ ;
       end if
      else
             Set $b=p_{2}$ ;
       end if
end for
Return $a$ if $a\cdot q_{a}\geq b\cdot q_{b}$ else return $b$
Algorithm 2 Ternary Search
Implementation details for Figure 5. The linear model satisfies for every $v\in[0,\infty)$,
$\overline{F}(v)=\min\left(\max(a-b\cdot v),0),1\right),$
where $b$ is uniformly sampled from $[1,5]$ and $a$ is uniformly sampled from $[1,b]$.
For the exponential demand we have for every $v\in[0,\infty)$,
$\overline{F}(v)=\min(\exp(a-b\cdot v),1),$
Finally, the stopping criterion used to derive the classical bound in Figure 5 satisfies for every desired level of maximin ratio $\epsilon>0$,
$N(\epsilon)=\frac{\log(\frac{\overline{v}-\underline{v}}{\overline{v}\cdot\epsilon})}{\log(3/2)}.$
@@ Appendix A Details on numerical evaluations in Section 3
Indeed, denote by $T_{N}=[a_{N},b_{N}]$, the interval obtained at iteration $N$ in Algorithm 2, by uni-modality of $\mathrm{Rev}\left(\cdot|\overline{F}\right)$ we know that $r_{F}$ is in $[a_{N},b_{N}]$, therefore we have:
$1-\frac{\mathrm{Rev}\left(a_{N}|\overline{F}\right)}{\mathrm{Rev}\left(r_{F}|\overline{F}\right)}$ $=\frac{r_{F}\cdot\overline{F}(r_{F})-a_{N}\cdot\overline{F}(a_{N})}{\mathrm{Rev}\left(r_{F}|\overline{F}\right)}$
$\leq\frac{(r_{F}-a_{N})\cdot\overline{F}(r_{F})}{\mathrm{Rev}\left(r_{F}|\overline{F}\right)}\leq\frac{b_{N}-a_{N}}{\mathrm{Rev}\left(r_{F}|\overline{F}\right)}=\frac{\overline{v}-\underline{v}}{\mathrm{Rev}\left(r_{F}|\overline{F}\right)}\left(\frac{2}{3}\right)^{N}\leq\frac{\overline{v}-\underline{v}}{\underline{v}}\left(\frac{2}{3}\right)^{N}$
@@
Therefore for $N\geq N(\epsilon)=\frac{\log(\frac{\overline{v}-\underline{v}}{\overline{v}\cdot\epsilon})}{\log(3/2)}$, we obtain the desired level of the maximin ratio.

### v2

@@ 6 Auxiliary Results
For any $r^{*}$ such that ${\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-empty, $\sup{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})=\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
@@
Step 1: We first show that, $\sup{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\leq\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Let $q\in{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ and let $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ such that ${r^{*}}$ is an optimal price for $F$ with associated conversion rate $q$. Consider $i\in\{0,\ldots,N\}$ such that ${r^{*}}\in[p_{i},p_{i+1})$ The crossing lemma implies that, $q=\overline{F}({r^{*}}-)\leq\min\left(\overline{G}_{\pazocal{F},{i-1}}({r^{*}}-),\overline{G}_{\pazocal{F},{i+1}}({r^{*}}-)\right)\leq\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}}),$ where the last inequality holds by definition of $\overline{U}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ when $\pazocal{C}$ is either $\pazocal{G}$ or $\pazocal{F}$.
@@
Step 2: Next, we show that $\sup{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\geq\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ by proving that $\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. To do so, we need to show that there exists a distribution $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ such that ${r^{*}}$ is an optimal price and $\overline{F}({r^{*}}-)=\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})$. We consider as a candidate the distribution $F_{\pazocal{C}}(\cdot|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}})$ defined in Equation 2.8. Let $\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+}=\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}}))$ and note that, $U_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$ and therefore this set is non-empty. Hence, by Lemma 3, $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$, which implies that, $F_{\pazocal{C}}(\cdot|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}})\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\subset\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
@@
To show that ${r^{*}}$ is the optimal price, we see from Lemma 4 that it suffices to show that ${r^{*}}\cdot\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})\geq\max_{i}p_{i}\cdot q_{i}$. Let $q\in{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ then by definition there exists $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ such that $\overline{F}({r^{*}}-)=q$ and ${r^{*}}\cdot\overline{F}({r^{*}}-)\geq p\cdot\overline{F}(p-)$ for every $p\in[\underline{v},\overline{v}]$. This implies that for every $i\in\{0,N+1\}$,
${r^{*}}\cdot\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})\stackrel{{\scriptstyle(a)}}{{\geq}}{r^{*}}\cdot q={r^{*}}\cdot\overline{F}({r^{*}}-)\geq p_{i}\cdot\overline{F}(p_{i}-)=p_{i}\cdot q_{i},$
@@
Lemma 8.
The ternary stopping criterion is a valid stopping criterion.
Proof of Lemma 8..
@@
Denote by $T_{N}=[a_{N},b_{N}]$, the interval obtained at iteration $N$ in Algorithm 1, by uni-modality of $\mathrm{Rev}\left(\cdot|F\right)$ we know that $r_{F}$ is in $[a_{N},b_{N}]$, therefore we have:
$1-\frac{\mathrm{Rev}\left(a_{N}|F\right)}{\mathrm{Rev}\left(r_{F}|F\right)}$ $=\frac{r_{F}\cdot\overline{F}(r_{F}-)-a_{N}\cdot\overline{F}(a_{N}-)}{\mathrm{Rev}\left(r_{F}|F\right)}$
$\leq\frac{(r_{F}-a_{N})\cdot\overline{F}(r_{F}-)}{\mathrm{Rev}\left(r_{F}|F\right)}\leq\frac{b_{N}-a_{N}}{\mathrm{Rev}\left(r_{F}|F\right)}=\frac{\overline{v}-\underline{v}}{\mathrm{Rev}\left(r_{F}|F\right)}\left(\frac{2}{3}\right)^{N}\leq\frac{\overline{v}-\underline{v}}{\underline{v}}\left(\frac{2}{3}\right)^{N}$
@@
Therefore for $N\geq N(\epsilon)=\frac{\log(\frac{\overline{v}-\underline{v}}{\underline{v}\cdot\epsilon})}{\log(3/2)}$, we obtain the desired level of the maximin ratio. ∎